# Beyond Trend and Periodicity: Guiding Time Series Forecasting with Textual Cues

### v2

@@ Appendix B Channel-wise performance on Weather-Captioned Dataset
The difficulty in predicting each channel varies, therefore, we present channel-wise performance in Table 5. The results demonstrate that TGForecaster, with the aid of external textual climate reports, significantly enhances forecasting accuracy across all channels. Notably, the model achieves over a 60% performance improvement in channels such as atmospheric pressure (p (mbar)), relative humidity (rh(%)), and vapor pressure deficit (VPdef (mbar)), which typically cannot be predicted reliably using historical time series data alone. The integration of external text cues has led to groundbreaking improvements in forecasting these parameters.
@@
Another noteworthy observation is that while TGForecaster is capable of predicting rainfall—unlike models that default to predicting near-zero averages—the performance improvement in these channels is modest. This is because rainfall is relatively scarce in this dataset, leading to large losses when rain is inaccurately predicted at the wrong times. Conversely, predicting the average value results in a smaller overall loss. This tendency explains why other models often opt for the average, avoiding the complex task of learning rainfall patterns. Nevertheless, accurate rainfall forecasting remains crucial in meteorological applications, underscoring our commitment to enhancing predictive accuracy in this area. The same principle also applies to other channels.
@@
Channel TGForecaster FITS DLinear PatchTST iTransformer IMP.
@@
Appendix C Visualization Weather-Captioned Dataset
@@
We provide the full visualization as Fig. 5. The TGForecaster shows great performance across all the channels. Even very hard ones such as Wind dir. It can also model the time series that totally independent with the weather such as the CO2 channel.
@@
Figure 5: Full visualization all channels on the 15000th test sample of Weather-Caption-Medium dataset. Blue line for ground truth, Red line for TGForecaster, Green line for PatchTST and Orange line for TGForecaster with swapping the news on the second and forth forecasting day.
Appendix D Weather Results w. w/o. RIN
@@
We compared the performance of models with and without Reversible Instance Normalization (RIN) on the weather-captioned-medium dataset, focusing on a 720-hour forecasting horizon. The model with RIN enabled achieved an MSE of 0.3428, whereas the model without RIN achieved a lower MSE of 0.2814. Results visualized in Fig. 6 show that the RIN-enabled model exhibits significant biases in many channels, particularly those with gradual trend shifts. This occurs because RIN removes the bias term from all instances, leaving the model unable to recognize relative bias and trend values. For instance, with RIN, temperature patterns in winter and summer are treated similarly, ignoring the typically higher and more variable temperatures in summer. Additionally, we noted pronounced shifting behavior coinciding with changes in captions, suggesting that the absence of bias information leads the model to over-rely on textual prompts, compensating for the missing data.
@@
Figure 6: Full visualization all channels on the 15000th test sample of Weather-Caption-Medium dataset. Blue line for ground truth, Red line for TGForecaster without RIN, Green line for TGForecaster with RIN enabled.
Appendix E Attention Map Visualization on Weather-Captioned
@@
We further visualize two cross-attention blocks to further investigate the TGForecaster. You are strongly advised to check the Tab. 1, Appendix K.3 and Fig. 5 while reading this part.
@@
In the second layer, however, there is a notable shift in attention dynamics. All channels, particularly channel 0, show intense focus on the sixth sentence. According to the channel definitions in Appendix K.3, channel 0 directly corresponds to atmospheric pressure. Channels 10 and 20, which are related to air density and CO2 concentration respectively—factors closely associated with pressure—also display relatively high attention scores. This suggests that the TGForecaster is capable of discerning the underlying relationships among the channels.
@@
(a) Layer1
(b) Layer2
(c) Layer3
Figure 7: Attention map of the "text guided channel independent" cross attention block on weather-captioned dataset, on on the 15000th test sample of Weather-Caption-Medium dataset. We use three cross attention block. The vertical axis stand for channels and horizon stand for the 7 sentences of the weather report summary.
Figure 8 presents the attention map of the modality mixer layer cross attention block in the weather-captioned dataset. The map, averaged across three cross attention layers, illustrates distinct patterns of attention for each channel. This diversity underscores the TGForecaster’s ability to adaptively extract time series embeddings tailored to the unique distribution characteristics of each channel, facilitated by textual inputs.
Notably, the channels for SWDR, PAR, and max.PAR display clear periodic patterns in their attention maps, aligning with observations from waveform visualizations. These patterns suggest that the TGForecaster effectively captures and utilizes periodic information from these environmental variables.
Furthermore, the channels labeled rain and raining show a particularly interesting behavior; they assign significantly higher attention scores to the exact time periods of rainfall within the look-back window. This behavior indicates that the TGForecaster is adept at identifying and prioritizing crucial temporal events specific to each channel, further enhancing its forecasting accuracy by focusing on relevant patterns where needed. This level of detail in attention allocation demonstrates the model’s capability to integrate contextual cues from textual data and further guide the time series forecasting.
Figure 8: Attention map of the modality mixer layer cross attention block on weather-captioned dataset, onon the 15000th test sample of Weather-Caption-Medium dataset. The attention map is averaged across three cross attention layers. We plot the attention map for each channel. The vertical axis stand for output time series patches and the horizon stand for the input time series patches embedding from PatchTST backbone.
Appendix F Comparison with More Baselines on Electricity-Captioned
We further compare with more baselines on Electricity-Captioned, including Autoformer, Fedformer, Informer, FiLM and TimesNet (Wu et al., 2021; Zhou et al., 2022a, 2021, b; Wu et al., 2023). TGForecaster shows dominant superior performance across these baselines, as shown in Tab. 6.
@@ Appendix F Comparison with More Baselines on Electricity-Captioned
Pred. Len. TGForecaster TGForecaster_120 Autoformer Fedformer Informer FiLM TimesNet
@@
Elec 720 0.193 0.209 0.254 0.244 0.373 0.236 0.22
Appendix G Full Result on Steam-100 Dataset
We show the comparison on Steam-100 Dataset in Tab. 7, and Tab, 8 with PatchTST. We use de-normed MAE as metric since the base volume of players of each game varies drastically, using normed metrics can lead to unfair comparison. The pretrain indicate that the model is jointly trained on all the games and each game is labeled by the channel discription. The Gnorm indicate that we apply the global normalization to preserve the player variation mentioned before. But it seems bring limited boost.
Table 7: Full result on Steam-100 dataset in de-normalized MAE. The best result is shown in green shaded bold font. The ones with performance boost over 10% is marked in red.
game_id TGTSF_pretrain
TGTSF_pretrain
@@ Appendix G Full Result on Steam-100 Dataset
Table 8: Cont. Full result on Steam-100 dataset in de-normalized MAE. The best result is shown in green shaded bold font. The ones with performance boost over 10% is marked in red.
game_id TGTSF_pretrain
TGTSF_pretrain
_Gnorm
TGTSF PatchTST IMP/%
@@
Appendix H Error Bar & Critical Difference Diagram
@@
As Tab. 9 indicate, TGForecaster shows stable performance across the benchmark. Even with extreme condition, it still maintains superior performance. It worth note that, we thought the relative large variance on weather dataset is caused by the different combination of the text description. But the FITS also shows large variance on this dataset which indicate it is hard to converge on this dataset.
We generate the critical difference plot on our result of four datasets (toy, Electricity, Weather-Medium, Weather-Large) with the default alpha as 0.05 as shown in Fig. 9. TGForecaster’s placement at the top of the critical difference plot, without intersecting with other lines, demonstrates its consistent and superior performance in terms of MSE compared to the other models. It indicates that with the help of external textual information, TGForecaster can handle complicated datasets.
Figure 9: The Critical Difference Plot on the TGForecaster and other baselines with alpha=0.05.
Table 9: The mean and std of TGForecaster on the three dataset in metrics of MSE.
Datasets TGForecaster FITS
@@
Appendix I Toy Dataset Details
I.1 Data source
@@
I.2 Dataset Structure
Toy/
|-- caption_emb/   // Pre-embeddings
|   |-- timestamp1.npy
|   |-- timestamp2.npy
|-- channel_1.npy     // channel description embedding data
|-- toy.csv     // time series data
Appendix J Electricity-Caption Details
J.1 Data source
@@ I.2 Dataset Structure
J.2 Dataset Structure
Elec-Captioned/
|-- caption_emb/   // Pre-embeddings
|   |-- Monday.npy
|   |-- Tuesday.npy
|   |-- Channel_1.npy
|   |-- Channel_2.npy
|-- electricity.csv     // time series data
Appendix K Weather-Caption Details
K.1 Data source

### v3

@@ Appendix G Channel-wise performance on Atmospheric Physics Dataset
The difficulty in predicting each channel varies, therefore, we present channel-wise performance in Table 5. The results demonstrate that FIATS, with the aid of external textual climate reports, significantly enhances forecasting accuracy across all channels. Notably, the model achieves over a 60% performance improvement in channels such as atmospheric pressure (p (mbar)), relative humidity (rh(%)), and vapor pressure deficit (VPdef (mbar)), which typically cannot be predicted reliably using historical time series data alone. The integration of external text cues has led to groundbreaking improvements in forecasting these parameters.
@@
Another noteworthy observation is that while FIATS is capable of predicting rainfall—unlike models that default to predicting near-zero averages—the performance improvement in these channels is modest. This is because rainfall is relatively scarce in this dataset, leading to large losses when rain is inaccurately predicted at the wrong times. Conversely, predicting the average value results in a smaller overall loss. This tendency explains why other models often opt for the average, avoiding the complex task of learning rainfall patterns. Nevertheless, accurate rainfall forecasting remains crucial in meteorological applications, underscoring our commitment to enhancing predictive accuracy in this area. The same principle also applies to other channels.
@@
Channel FIATS FITS DLinear PatchTST iTransformer IMP.
@@
Appendix H Performance Visualization on Atmospheric Physics Dataset
@@
We provide the full visualization as Fig. 8. The FIATS shows great performance across all the channels. Even very hard ones such as Wind dir. It can also model the time series that totally independent with the weather such as the CO2 channel.
@@
Figure 8: Full visualization all channels on the 15000th test sample of Weather-Caption-Medium dataset. Blue line for ground truth,Red line for FIATS, Green line for PatchTST and Orange line for FIATS with swapping the news on the second and forth forecasting day.
Appendix I Weather Results w. w/o. RIN
@@
We compared the performance of models with and without Reversible Instance Normalization (RIN) on the Atmospheric Physics-medium dataset, focusing on a 720-hour forecasting horizon. The model with RIN enabled achieved an MSE of 0.3428, whereas the model without RIN achieved a lower MSE of 0.2814. Results visualized in Fig. 9 show that the RIN-enabled model exhibits significant biases in many channels, particularly those with gradual trend shifts. This occurs because RIN removes the bias term from all instances, leaving the model unable to recognize relative bias and trend values. For instance, with RIN, temperature patterns in winter and summer are treated similarly, ignoring the typically higher and more variable temperatures in summer. Additionally, we noted pronounced shifting behavior coinciding with changes in captions, suggesting that the absence of bias information leads the model to over-rely on textual prompts, compensating for the missing data.
@@
Figure 9: Full visualization all channels on the 15000th test sample of Weather-Caption-Medium dataset. Blue line for ground truth, Red line for FIATS without RIN, Green line for FIATS with RIN enabled.
Appendix J Ablation Study on Causal Relationship Extraction
FIATS is designed to learn causal relationships between events described in text and their corresponding time series patterns. While not explicitly an alignment model, it effectively aligns the semantic meaning of text with the time series data it impacts. The model generates time series patterns guided by the textual information, and its performance varies based on the quality of the text input:
  1. 1.
    Training with Meaningful and Relevant Text:
    • Inference with Similar Text: Produces strong results by accurately extracting causal relationships between events in the text and time series patterns.
  2. 2.
    Training with Zero/Random Text:
    • Inference with Any Text: Produces results equivalent to PatchTST, as no additional information is present in the text. The model relies solely on the time series data, ignoring the random text.
  3. 3.
    Training with Meaningful Text, Inference with Incorrect Text:
    • Inference with Incorrect Text: Results are poor, as the model relies on the misleading text input and generates patterns based on incorrect or irrelevant information.
We detail the FIATS performance under different text conditions in the following table:
Table 6: FIATS performance under different training and testing conditions. We report the result of forecasting horizon 96 on Atmospheric Physics-medium dataset using MiniLM embedding.
Train with
Good Zero Random
Test with Good
0.186
(captures causal relationships)
0.249
(corrupted random patterns)
0.251
(similar to PatchTST)
Zero
0.724
(corrupted repetive patterns)
0.249
(similar to PatchTST)
0.254
(similar to PatchTST)
Random
0.615
(corrupted random patterns)
0.249
(similar to PatchTST)
0.250
(similar to PatchTST)
The results of the ablation study provide strong evidence that FIATS relies on capturing causal relationships between time series patterns and dynamic news, rather than simply treating text as auxiliary input. When trained with meaningful and correlated news, the model demonstrates its ability to effectively extract these relationships, yielding strong predictive performance. This highlights FIATS’s capacity to align the semantic meaning of text with time series patterns in a causally meaningful way.
On the other hand, when trained with good text but tested with random or misleading text, the model produces poor predictions because it continues to rely on the input text, even when it is inaccurate or irrelevant. This further underscores the model’s dependence on the quality of the textual input rather than merely defaulting to learned time series patterns.
Interestingly, when trained with bad or random text, FIATS fails to establish causal relationships and instead reverts to PatchTST-level performance, indicating it falls back to relying solely on time series data. Furthermore, when subsequently tested with good text, the model trained on bad text still ignores the input entirely, suggesting it stops depending on textual input when the training data lacks meaningful causal relationships.
These results collectively demonstrate that FIATS’s strength lies in its ability to extract and leverage causal relationships between text and time series data. The model’s performance is tightly coupled with the quality and relevance of the textual input, validating the centrality of causal alignment in its design and functionality.
These outcomes demonstrate that IATSF effectively achieves alignment in the “event” space, linking events described in the text to the corresponding time series patterns.
Appendix K Attention Map Visualization on Atmospheric Physics
@@
We further visualize two cross-attention blocks to further investigate the FIATS. You are strongly advised to check the Tab. 13, Appendix P.4.5 and Fig. 8 while reading this part.
@@
In the second layer, however, there is a notable shift in attention dynamics. All channels, particularly channel 0, show intense focus on the sixth sentence. According to the channel definitions in Appendix P.4.5, channel 0 directly corresponds to atmospheric pressure. Channels 10 and 20, which are related to air density and CO2 concentration respectively—factors closely associated with pressure—also display relatively high attention scores. This suggests that the FIATS is capable of discerning the underlying relationships among the channels.
@@
(a) CASM Layer1
(b) CASM Layer2
(c) CASM Layer3
Figure 10: Attention map of the "CASM" cross attention block on the 15000th test sample of Atmospheric Physics dataset dataset. We use three cross attention block. The vertical axis stand for channels and horizon stand for the 7 sentences of the weather report summary.
Figure 11 presents the attention map of the modality mixer layer cross attention block in the Atmospheric Physics dataset. The map, averaged across three cross attention layers, illustrates distinct patterns of attention for each channel. This diversity underscores the FIATS’s ability to adaptively extract time series embeddings tailored to the unique distribution characteristics of each channel, facilitated by textual inputs.
Notably, the channels for SWDR, PAR, and max.PAR display clear periodic patterns in their attention maps, aligning with observations from waveform visualizations. These patterns suggest that the FIATS effectively captures and utilizes periodic information from these environmental variables.
Furthermore, the channels labeled rain and raining show a particularly interesting behavior; they assign significantly higher attention scores to the exact time periods of rainfall within the look-back window. This behavior indicates that the FIATS is adept at identifying and prioritizing crucial temporal events specific to each channel, further enhancing its forecasting accuracy by focusing on relevant patterns where needed. This level of detail in attention allocation demonstrates the model’s capability to integrate contextual cues from textual data and further guide the time series forecasting.
Figure 11: Attention map of the modality mixer layer cross attention block on Atmospheric Physics dataset, on the 15000th test sample of Weather-Caption-Medium dataset. The attention map is averaged across three cross attention layers. We plot the attention map for each channel. The vertical axis stand for output time series patches and the horizon stand for the input time series patches embedding from PatchTST backbone.
Appendix L Comparison with More Baselines on Electrical Utility
We further compare with more baselines on Electrical Utility, including Autoformer, Fedformer, Informer, FiLM and TimesNet [30, 31, 16, 32, 33]. FIATS shows dominant superior performance across these baselines, as shown in Tab. 7.
@@
Pred. Len. FIATS FIATS_120 Autoformer Fedformer Informer FiLM TimesNet
@@
Elec 720 0.193 0.200 0.254 0.244 0.373 0.236 0.22
Appendix M Full Result on GAUD Dataset
We show the comparison on GAUD Dataset in Tab. 8, and Tab, 9 with PatchTST. We use de-normed MAE as metric since the base volume of players of each game varies drastically, using normed metrics can lead to unfair comparison. The pretrain indicate that the model is jointly trained on all the games and each game is labeled by the channel discription. The Gnorm indicate that we apply the global normalization to preserve the player variation mentioned before. But it seems bring limited boost.
Table 8: Full result on GAUD dataset in de-normalized MAE. The best result is shown in green shaded bold font. The ones with performance boost over 10% is marked in red.
game_id IATSF_pretrain
IATSF_pretrain
@@
Table 9: Cont. Full result on GAUD dataset in de-normalized MAE. The best result is shown in green shaded bold font. The ones with performance boost over 10% is marked in red.
game_id IATSF_pretrain
IATSF_pretrain
_Gnorm
IATSF PatchTST IMP/%
@@
Appendix N Error Bar & Critical Difference Diagram
@@
As Tab. 10 indicate, FIATS shows stable performance across the benchmark. Even with extreme condition, it still maintains superior performance. It worth note that, we thought the relative large variance on weather dataset is caused by the different combination of the text description. But the FITS also shows large variance on this dataset which indicate it is hard to converge on this dataset.
We generate the critical difference plot on our result of four datasets (toy, Electricity, Weather-Medium, Weather-Large) with the default alpha as 0.05 as shown in Fig. 12. FIATS’s placement at the top of the critical difference plot, without intersecting with other lines, demonstrates its consistent and superior performance in terms of MSE compared to the other models. It indicates that with the help of external textual information, FIATS can handle complicated datasets.
Figure 12: The Critical Difference Plot on the FIATS and other baselines with alpha=0.05.
Table 10: The mean and std of FIATS on the three dataset in metrics of MSE.
Datasets FIATS FITS
@@
Appendix O Experiment on Time-MMD
Initially, we planned to include the Time-MMD dataset as a real-world scenario to benchmark our method. However, upon evaluation, we found that this dataset suffers from significant flaws and poor organization. As a result, we have decided to include it in the appendix as supplementary material and highlight some of its issues.
O.1 Results on Time-MMD
Despite the dataset’s limitations, our FIATS demonstrates state-of-the-art performance on the Time-MMD dataset, as shown in Tab. 11. In several subsets, FIATS achieves a performance improvement exceeding $50\%$, showcasing its robustness and effectiveness even when applied to a flawed dataset.
Table 11: Results on TimeMMD.
Agriculture Climate
FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG
0.07 0.09 0.17 0.24 1.02 1.24
0.11 0.12 0.19 0.34 1.02 1.24
0.16 0.15 0.24 0.43 1.03 1.24
0.17 0.18 0.29 0.51 1.03 1.24
Economy Traffic
FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG
0.14 0.10 0.35 0.14 0.188 0.24
0.16 0.11 0.4 0.15 0.19 0.25
0.17 0.14 0.41 0.17 0.19 0.24
0.2 0.14 0.37 0.18 0.24 0.29
Socialgood Security
FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG
0.66 0.82 0.87 68.51 112.76 118.49
0.75 0.93 0.99 85.81 115.33 119.09
0.78 1.01 1.07 89.26 117.19 121.08
0.86 1.05 1.13 92.89 118.03 123
Energy Health
FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG
0.11 0.14 0.16 1.38 1.12 1.55
0.21 0.24 0.27 1.82 1.4 1.88
0.3 0.32 0.35 2.01 1.48 1.91
0.36 0.44 0.46 3.04 1.53 1.97
Environment
FIATS Time-MMD-multi-AVG Time-MMD-uni-AVG
0.32 0.32 0.35
0.34 0.35 0.38
0.35 0.37 0.47
0.37 0.41 0.4
O.2 Poor Data Quality of Time-MMD
Unbalanced Dataset.
Many subsets of the Time-MMD dataset span extended periods to collect as many valid numerical data points as possible, with some dating back to the 1980s. However, textual information from earlier periods is largely absent, resulting in a lack of corresponding records for these time intervals. In contrast, recent years have seen a surge in text articles available online. This imbalance creates challenges for the model: during training, it cannot effectively learn correlations between text and time series due to sparse or missing text data, while during inference, the model is inundated with abundant textual information. As a result, the dataset becomes inherently unbalanced.
Meaningless Placeholder Text
Moreover, some entries in the dataset include placeholder text generated by large language models, indicating their inability to produce meaningful output due to insufficient input data. For instance, in the "Agriculture_search.csv" file, over $80\%$ of the entries consist of statements like: "Since there is no relevant information, I am unable to provide any objective facts, insights, analysis, or predictions about the United States broiler market. This search result is not relevant to United States Retail Broiler or Retail Chicken. It appears to be an advertisement for a perfume and has no connection to the topic."
Similarly, in the "Economy_search.csv", other entries state: "After reviewing the search results, I found that most of the information is not relevant to making predictions about the Economy. However, I was able to extract some useful information, which I have summarized below: NA." While these entries appear to be valid text data, they offer no meaningful or actionable information, further compounding the issue of data imbalance.
These meaningless information are all over the whole dataset, making the text validity of this dataset doubtful.
Information Leakage.
Another significant issue with the dataset is information leakage. The dataset creators used large language models to process reports or search results and generate "fact" and "prediction" entries. However, in some cases, the reports or search results directly contain the actual values to be predicted, leading to severe information leakage.
For example, in the "Agriculture_report.csv" file, the "fact" entry for the date 2019-02-04 states: "The National Composite Weighted Average for 1/31/19 is 92.04 compared to 94.22 a week earlier, and 91.66 a year ago." This directly provides the target value to be predicted. Such instances of information leakage are pervasive throughout the dataset and significantly compromise the reliability of the results derived from it.
Appendix P IATSF Benchmark Datasets
We designed the IATSF benchmark to include four datasets of varying complexity, each tailored to evaluate specific aspects of model performance. Together, these datasets form a progression from simple, interpretable scenarios to challenging, real-world applications, providing a comprehensive evaluation framework for text-guided time series forecasting models.
1. Toy Dataset The Toy dataset is intentionally designed with simple and straightforward patterns, making it easy to analyze and interpret. However, the dataset includes sudden changes in patterns that are impossible to predict without text guidance. This ensures the model’s ability to adhere to textual cues is effectively tested in a controlled environment. It serves as a foundation for validating whether the model can extract and use textual guidance to forecast time series.
2. Electricity Dataset The Electricity dataset introduces real-world data with common textual features like day of the week or public holidays. While the textual information is relatively simple, it tests the model’s ability to utilize such structured cues for forecasting. Additionally, as a widely-used off-the-shelf dataset, it allows for easy comparison with existing methods, providing a baseline for evaluating IATSF’s performance.
3. Atmospheric Physics Dataset The Atmospheric Physics dataset represents a semi-controlled environment designed to rigorously test the model’s ability to learn causal relationships between text and time series patterns. It also evaluates the model’s text-guided channel independence and generalizability. By simulating a scenario where text and time series data are strongly correlated, this dataset bridges the gap between controlled tests and more complex real-world challenges.
4. GAUD Dataset The GAUD dataset is a fully real-world dataset that tests IATSF in a practical industrial context. Its patterns are noisy and random, making it highly challenging. This dataset showcases IATSF’s ability to perform well in realistic scenarios.
Comprehensive Benchmark Objectives
The IATSF benchmark is designed to address multiple objectives:
  • Interpretability and Validation: The simpler Toy and Electricity datasets help researchers validate their models and understand their behavior in controlled environments.
  • Performance Testing in Complex Scenarios: The Atmospheric Physics and GAUD datasets challenge the models in semi-controlled and real-world settings, ensuring they are robust and capable of handling practical applications.
This benchmark is not merely a ranking tool but a framework to help researchers analyze and improve their models’ behaviors across varying levels of complexity. We see this as a starting point for the community and hope it will inspire researchers to contribute additional datasets, further expanding and enriching the IATSF benchmark for future advancements in this field.
P.1 Metadata for Datasets
We show the metadata for IATSF Datasets in Table 12.
Table 12: Datasets Metadata
Dataset Length Time span TS Sampling Rate # of Channels # of Dynamic News each step Textual update rate Notes
Toy 300,000 N/A N/A 1 1$\sim$3 Every Step Sinusoidal wave with a single channel
Electrical Utility 26,304 2011-01-01 to 2015-12-31 1 hour 321 1$\sim$3 Daily Just the Electricity Dataset
Atmospheric Physics 525,600 2014-01-01 to 2023-12-31 10 minutes 21 7 Every 6 hours Weather data with 21 channels. Three set of textual cues for combination.
GAUD Varies 2005 to 2024 1 Day 1 (each game) Varies Varies Each game has historical data from its prelaunch to 2024.
P.2 Toy Dataset Details
@@
P.3 Electricity-Caption Details
@@
P.4 Atmospheric Physics Details
P.4.1 Data source